OpenStudy (anonymous):

Isolate y in the equation 3x + 2y = 8

OpenStudy (anonymous):

3x +2y=8 2y=.3x+8 y=3/2x+4 is the answer :)

OpenStudy (anonymous):

y=-3/2x+8 sorri:/

OpenStudy (anonymous):

y=-3/2x+4

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hope that helps:)